OpenStudy (anonymous):

^20sqrtx12 use rational exponents to simplify the radical. if rational exponents appear after simplifying, write the answer in radical notation. a. 4sqrtx3 b. 5sqrtx3 c. 5sqrtx d. 4sqrtx

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt[20]{x^{12}}\] \[\Large x^{\frac{12}{20}}\] \[\Large x^{\frac{3}{5}}\] \[\Large \sqrt[5]{x^{3}}\] So \[\Large \sqrt[20]{x^{12}}=\sqrt[5]{x^{3}}\]